# Human-in-the-Loop Optimization with Model-Informed Priors

### v1

@@ 6. Discussion
Optimization plays a key role in HCI, particularly for adaptive user interfaces. Prior work tends to follow one of two distinct paradigms: On one hand, model-based approaches leverage established user models to compute optimal designs, but they ignore real-time observations, unable to truly personalize to individual users or dynamic contexts. On the other hand, human-in-the-loop optimization adapts directly to individuals through online evaluation, but often requires significant exploration, making it slow to converge. In this work, we propose a novel concept, utilizing the advantages of both: pre-training optimizers using established user models to accelerate adaptation in human-in-the-loop settings. We define a novel conceptual framework, HOMI, which unifies model-based and interactive optimization by using synthetic users to train an adaptive optimizer prior to deployment. We provide the core components and stages of this framework and instantiate it through a concrete method: NAF+. NAF+ is a new optimization method that integrates a deep-learning-based acquisition function, trained via reinforcement learning on synthetic users to support efficient real-time adaptation. To extend its applicability to more realistic HCI scenarios, we build upon the Neural Acquisition Function (NAF) and introduce two key innovations: (1) the ability to handle dynamic trade-offs between multiple design objectives via weighted inputs, and (2) a novelty-aware fallback mechanism that detects novel users and adapts accordingly. Together, these contributions enable NAF+ to achieve fast, personalized adaptation in realistic, multi-objective design tasks. Our synthetic tests and user study validate the efficacy of our approach. Despite these promising results, several open questions and limitations remain.
@@
Traditional model-based personalization requires fitting a parameterized user model to each new individual, which often involves complex and error-prone user inference pipelines before any adaptation can occur (Beaumont et al., 2002). Although recent learning-based inference approaches (e.g., amortized inference (Moon et al., 2023, 2024)) have been proposed to mitigate this burden, model inference remains a challenging step to support real-time adaptation to the diverse user characteristics. In contrast, HOMI adopts a different philosophy: we do not seek to model any single user perfectly. Instead, we train the optimizer on a broad distribution of synthetic users, enabling it to learn adaptation strategies that generalize across populations. At deployment, no explicit user modeling or inference is needed; the system simply adapts based on observed performance, guided by the meta-learned policy. This shift—away from high-fidelity modeling toward population-level diversity and meta-adaptation offers a more scalable and robust path for real-time personalization.
@@
Traditionally, user models have been developed to characterize and understand human performance, but rarely to serve a generative role. As a result, there has been little incentive to create or carefully fit such models for many interaction modalities, which explains why reliable parameter estimates are often missing, even when suitable models are available. This creates a fundamental challenge for our workflow: HOMI works best when models and parameters exist which allows our workflow to skip the parameter fitting step, yet in most realistic cases today, parameter fitting is likely to be needed. Our HOMI and NAF+ provide a new purpose for user models: shifting them from passive tools for understanding or characterizing human behavior, to the enabler of active interface optimization and adaptation. We hope this re-positioning will encourage the community to not only develop user models but also to document and share parameter values across different interactions and modalities, ultimately enabling broader deployment of synthetic-user-driven optimization.
@@
Our approach is built upon the broad framework of BO and RL, which have shown strong scalability and generalizability across applications. In future work we will apply the HOMI framework to other interactive systems where well-established performance or cognitive models exist. Promising examples include menu optimization, gaze-based interfaces, and gesture input—scenarios where task-specific models can be leveraged to generate meaningful synthetic users and enable pretraining. We can then further validate the flexibility of our approach and demonstrate its potential as a general solution for scalable, adaptive interface optimization.
@@
HOMI is grounded in the use of parameterized user models (e.g., Fitts’ law) to simulate synthetic users and enable large-scale pretraining. While this provides a scalable and principled way to approximate diverse user behaviors, it requires well-established models for a given interaction task. Such models are limited, and moreover, they are usually not available for more complex interactions involving rich cognitive processes or sequential decision-making or motions. Recent advances in Computational Rationality and reinforcement learning–based modeling of cognitive and motor processes  (Oulasvirta et al., 2022; Murray-Smith et al., 2022) offer promising alternatives for simulating realistic human behavior in such scenarios. Additionally, generative models, such as variational autoencoders, diffusion models, or large language models, are increasingly used to model human-like behavior from data, and could be adapted to synthesize user interaction patterns in tasks where explicit models are lacking (Li et al., 2024; Park, 2025). Future work could incorporate these approaches to expand the scope of HOMI, or explore hybrid user simulation strategies that blend analytical models with data-driven user representations. These directions can significantly broaden the applicability of our framework across a wider spectrum of HCI domains.
@@ 7. Conclusion
We introduce HOMI, a novel framework for HILO that leverages model-informed pre-training to accelerate real-time interface adaptation. By synthesizing large populations of simulated users from established user models, HOMI enables the training of powerful optimizers prior to deployment, bridging the gap between traditional model-based design and purely interactive approaches. To demonstrate the framework, we proposed NAF+, a concrete method that integrates a deep reinforcement learning–based acquisition function, dynamic multi-objective adaptation, and a novelty-aware fallback mechanism. Together, these components enable efficient and robust optimization across diverse users and changing objectives. Our results suggest that this new concept and approach jointly offer a scalable and generalizable path for future personalization systems: one that moves beyond per-user modeling and instead learns population-level adaptation strategies that generalize in deployment. We believe this work opens up a new direction for optimization in HCI, one that treats user models not just as analytical tools, but as generative engines for training adaptive, intelligent design systems. Finally, we hope this work provides a new lens for bridging in vitro (in simulation) with in situ (with real users) optimization, enabling more streamlined and adaptive future interfaces.
@@ 1. NAF+
We set the naf as a six-layer fully connected neural network, with each hidden layer containing 256 nodes and Tanh as the activation function. The naf is trained using Proximal Policy Optimization (PPO) (Schulman et al., 2017), where the critic network shares the same architecture as the actor (i.e., naf). The actor and critic learning rates are set to 0.0003 and 0.001, respectively. PPO is run for $K=10$ epochs with a clipping parameter of 0.15. To encourage generalization across varying objective weightings, the objective weights are randomly sampled during training. To be more specific, the weight on the first Sphere function, $\alpha$, is drawn uniformly from $[0,1]$, and the weight on the second is set to $1-\alpha$. This two-dimensional weight vector $(\alpha,1-\alpha)$ is also included as part of the input to the naf, allowing the acquisition function to adapt its strategy based on the current configuration. Note, when deploying on the test functions, the weights are given to NAF+ as a part of the input. We denote this method as NAF+ in the rest of this test.
@@ 5. Standard Bayesian optimization (BO)
The final condition is standard Bayesian optimization (BO). We used Expected Improvement as the acquisition function and a typical Single-Task GP implemented in BoTorch library as the surrogate model [^4] https://botorch.org/. We set BO with 6 random explorations followed by 14 optimization steps.
@@
Test 1.1 is designed to specifically evaluate the value of weight information. To this end, we sampled 20 test functions from the exact same range as the one used for training. Since the test functions are fully overlap with the training range, we exclude the condition with novel detection and additional EI (NAF-w/o-EI) from this test The log regret results are presented in Figure 6 left panel. Overall, we can observe that NAF+ started from an overall better performance and constantly outperformed other meta-BO approaches, such as NAF-w/o-Weight and TAF. Also, while TAF showed promising performance (most regret is less than $10^{-1}$), it still performed worse than NAF+ and NAF-w/o-Weight. The primary reason is that NAF+ is trained by nearly unlimited number of synthetic agents, allowing it to fully generalize across new tasks. On the other hand, TAF is based on a limited set of prior synthetic users. Lastly, within the first 10 iterations, all meta-BO approaches demonstrated better performance than standard BO, which starts from scratch. This highlights the benefit of prior experience.
Figure 6. The results of Test 1.1 (left) and Test 1.2 (right). Note that the value plot here is the $log_{10}$ of the regret.
@@ C.2. Configurations of each Condition
Similar to the previous test, Test 2.1 is designed to evaluate the value of the weight information. We sampled 20 test functions from the exact same range as the one used for training. We again exclude the condition with novel detection (NAF-w/o-Weight) from this test as there will not be novel functions. The log regret results are presented in Figure 7, left panel. NAF+ and NAF-w/o-Weight start from an overall better performance than the others. Especially, NAF+ converge to optimal performance around the second iteration, and NAF-w/o-Weight converges around the fifth iteration, showing NAF+ still benefits from the additional weight information as a part of the input. We observed that TAF with 15 prior models struggled to improve significantly in this case. This might be because of the higher number of parameters (2 in Test 1 and 6 in Test 2). To cover this number of dimensions, potentially more models will be required. Finally, same as the previous test, Standard BO starts from a worse performance and gradually converges toward optimal performance.
@@
Test 2.2 evaluates the value of merging EI and the novelty detector. To create a scenario where the new user is completely outside of the training range, we significantly shift the range of sampling the model parameters. The mean for sampling synthetic users are set as $[180,120,0.0145,1.6,0.03,2.0]$ for $a,b,\alpha_{x},\sigma_{\alpha_{x}},\alpha_{y},\sigma_{\alpha_{y}}]$, respectively, and the standard deviation remains the same as. These ranges are significantly different from the parameter settings for training, ensuring all sampled users are novel. We sampled 20 test functions from the test range. The log regret results are presented in Figure 7, right panel. We find that NAF+ and NAF-w/o-EI start from overall better performance, but NAF-w/o-EI does not improve significantly throughout the iterations, while NAF+ continually improves. This, again, highlights the potential issue of not incorporating an additional EI to handle out-of-distribution cases: naf can not generalize to other cases once it is fully trained to a certain range. TAF and standard BO perform well in this task. However, they start from a worse performance. BO, especially, requires 4 to 5 iterations to catch up with other conditions.
@@ C.5. Summary of Test 2
In summary, both Test 1 (a controlled synthetic benchmark) and Test 2 (a more realistic simulation grounded in HCI literature) converge on the same conclusion: NAF+, through its use of explicit weight input and novelty-aware EI, achieves greater sample efficiency and robustness across a wide range of user profiles. The positive outcomes of Test 2 further motivate our user study, which targets typing performance on mid-air keyboards, highly relevant with Test 2 – a touch-based keyboard for mobile devices.
Figure 7. The results of Test 2.1 (left) and Test 2.2 (right). Note that the value plot here is the $log_{10}$ of the regret.

### v2

@@ 6. Discussion
We propose, HOMI, a novel concept which pre-trains meta-BO using user models to accelerate adaptation in human-in-the-loop settings. We further introduce NAF+, a new optimization method that integrates a deep-learning-based acquisition function, trained via reinforcement learning on synthetic users to support efficient real-time adaptation. Our synthetic tests and user study validate the efficacy of our approach. Despite these promising results, several open questions and limitations remain.
@@
Traditional model-based personalization requires fitting a parameterized user model to each user, which often involves complex model inverse steps before any adaptation can occur (Beaumont et al., 2002). Although recent learning-based inference approaches (e.g., amortized inference (Moon et al., 2023, 2024)) have been proposed to mitigate this challenging, model inference remains a challenging step to support real-time adaptation to the diverse user characteristics. In contrast, HOMI adopts a different philosophy: we do not seek to model any single user perfectly. Instead, we train the optimizer on a broad distribution of synthetic users, enabling it to learn adaptation strategies that generalize across populations. At deployment, no explicit user modeling or inference is needed; the system simply adapts based on observed performance, guided by the meta-learned policy. This shift from high-fidelity modeling toward population-level diversity and meta-adaptation offers a more scalable and robust path for real-time personalization.
@@
User models in HCI have traditionally been developed as descriptive tools mainly for characterizing and predicting human performance, but not meant for interface generation and adaptation. Thus, even for well-studied behaviors such as pointing or typing, available models were rarely fitted comprehensively across devices or modalities, because there was little downstream use for such parameter values. This gap creates a potential challenge for HOMI: our workflow benefits most when reliable models and parameter ranges already exist, enabling us to bypass the parameter-fitting step. Our HOMI and NAF+ introduce a fundamentally different purpose for user models: transforming them from passive analytic tools into active enablers of interface optimization, adaptation, and even generation. This re-positioning creates an actionable incentive for the community: rather than treating user models as isolated theoretical artifacts, we encourage documenting, validating, and sharing their parameter values across interactions, devices, and modalities. A collaboratively maintained gallery of models and parameter distributions. That is, not just equations, but reusable parameter sets, would unlock scalable synthetic-user–driven optimization and significantly broaden the applicability of HOMI-style workflows across HCI.
@@
Our approach is built upon the broad framework of BO and RL, which have shown strong scalability and generalizability across applications. In future work, we will apply the HOMI framework to other interactive systems where well-established performance or cognitive models exist. Promising examples include menu optimization, gaze-based interfaces, gesture input, and general scenarios where task-specific models can be leveraged to generate meaningful synthetic users and enable pretraining. We can then further validate the flexibility of our approach and demonstrate its potential as a general solution for scalable, adaptive interface optimization.
@@
HOMI is grounded in the use of parameterized user models to simulate synthetic users and enable large-scale pretraining. While this provides a scalable and principled way to approximate diverse user behaviors, it requires established models for a given interaction task. Such models might not be available for more complex interactions that involve rich cognitive processes, sequential decision-making, or dexterous motions. Here, we see an opportunity for the use of more advanced modeling approaches that have recently emerged in HCI. Specifically, advances in Computational Rationality and reinforcement learning–based modeling of cognitive and motor processes  (Oulasvirta et al., 2022; Murray-Smith et al., 2022; Liao and Holz, 2025) offer promising alternatives for simulating realistic human behavior in such scenarios. When combined with biomechanical models that provide realistic musculoskeletal dynamics (Ikkala et al., 2022; Moon et al., 2024, 2024; Liao et al., 2022; Ikkala et al., 2022; Fischer et al., 2024), this approach offers a powerful path toward general-purpose human simulators. Unlike traditional analytical models that are tied to specific tasks, these reinforcement learning-based models can learn to reproduce a range of sensorimotor behaviors through learning. This opens the door to synthetic users that can generalize across tasks, contexts, and devices, further broadening the applicability of HOMI. Additionally, generative models, such as variational autoencoders, diffusion models, or large language models, are increasingly used to model human-like behavior from data, and could be adapted to synthesize user interaction patterns in tasks where explicit models are lacking (Li et al., 2024; Park, 2025). Future work could incorporate these approaches to expand the scope of HOMI, or explore hybrid user simulation strategies that blend analytical models with data-driven user representations. These directions can significantly broaden the applicability of our framework across a wider spectrum of HCI domains.
@@ 7. Conclusion
We introduce HOMI, a novel framework for HILO that leverages model-informed pre-training to accelerate real-time interface adaptation. By synthesizing large populations of simulated users from user models, HOMI enables the training of powerful optimizers prior to deployment, bridging the gap between traditional model-based design and purely interactive approaches. To demonstrate the framework, we proposed NAF+, a concrete method that integrates a deep reinforcement learning–based acquisition function, dynamic multi-objective adaptation, and a novelty-aware fallback mechanism. Together, these components enable efficient and robust optimization across diverse users and changing objectives. Our results suggest that this new concept and approach jointly offer a scalable and generalizable path for future personalization systems: one that moves beyond per-user modeling and instead learns population-level adaptation strategies that generalize in deployment. We believe this work opens up a new direction for optimization in HCI, one that treats user models not just as analytical tools, but as generative engines for training adaptive, intelligent design systems. Finally, we hope this work provides a new lens for bridging in vitro (in simulation) with in situ (with real users) optimization, enabling more streamlined and adaptive future interfaces.
8. Open Science
NAF+ is released at https://github.com/yichiliao/homi. We hope that our implementation can encourage other researchers to build other applications based on our proposed workflow and approach.
Acknowledgements.
The authors sincerely thank Aakar Gupta, Ruta Desai, Antti Oulasvirta, and Aleksi Ikkala for their insightful discussions. Yi-Chi Liao was partially supported by the ETH Zurich Postdoctoral Fellowship Programme. Hee-Seung Moon is supported by the National Research Foundation of Korea (RS-2025-00521470).
@@ 1. NAF+
We set the naf as a six-layer fully connected neural network, with each hidden layer containing 256 nodes and Tanh as the activation function. The naf is trained using Proximal Policy Optimization (PPO) (Schulman et al., 2017), where the critic network shares the same architecture as the actor (i.e., naf). The actor and critic learning rates are set to 0.0003 and 0.001, respectively. PPO is run for $K=10$ epochs with a clipping parameter of 0.15. To encourage generalization across varying objective weightings, the objective weights are randomly sampled during training. To be more specific, the weight on the first Sphere function, $\alpha$, is drawn uniformly from $[0,1]$, and the weight on the second is set to $1-\alpha$. This two-dimensional weight vector $(\alpha,1-\alpha)$ is also included as part of the input to the naf, allowing the acquisition function to adapt its strategy based on the current configuration. Note, when deploying on the test functions, the weights are given to NAF+ as a part of the input. Finally, the pre-defined novelty threshold $\tau$ is set to be 0.1. We denote this method as NAF+ in the rest of this test.
@@ 5. Standard Bayesian optimization (BO)
The final condition is standard Bayesian optimization (BO). We used Expected Improvement as the acquisition function and a typical Single-Task GP implemented in the BoTorch library as the surrogate model [^4] https://botorch.org/. We set BO with 6 random explorations followed by 14 optimization steps.
@@
Test 1.1 is designed to specifically evaluate the value of weight information. To this end, we sampled 20 test functions from the exact same range as the one used for training. Since the test functions are fully overlap with the training range, we exclude the condition with novel detection and additional EI (NAF-w/o-EI) from this test The log regret results are presented in Figure 8 left panel. Overall, we can observe that NAF+ started from an overall better performance and constantly outperformed other meta-BO approaches, such as NAF-w/o-Weight and TAF. Also, while TAF showed promising performance (most regret is less than $10^{-1}$), it still performed worse than NAF+ and NAF-w/o-Weight. The primary reason is that NAF+ is trained by a nearly unlimited number of synthetic agents, allowing it to fully generalize across new tasks. On the other hand, TAF is based on a limited set of prior synthetic users. In addition, within the first 10 iterations, all meta-BO approaches demonstrated better performance than standard BO, which starts from scratch. This highlights the benefit of prior experience. To further analyze the benefit of incorporating the weight information, we run independent-samples t-tests to specifically compare the resulting performance of NAF+ and NAF-w/o-Weight at each iteration. We found significant differences in iteration 5-20, all $p<0.05$, confirming that NAF overall outperforms NAF-w/o-Weight.
Figure 8. The results of Test 1 (double-sphere functions): Test 1.1 (evaluating weight information) is shown in left, and Test 1.2 (evaluating expected improvement) is shown in right. The value plot here is the $log_{10}$ of the regret. In Test 1.1, the * signs denote the iterations (5-20) where a significant difference between NAF+ and NAF-w/o-Weight was found. The results show that NAF+ benefits from the weight information and thus can achieve better performance than NAF-w/o-Weight by a more informed optimization strategy. Meanwhile, the novelty detector allows NAF+ to rely on its GP and EI to address tasks that are outside of the training distribution, whereas NAF-w/o-EI can not adapt to novel tasks.
@@ C.2. Configurations of each Condition
Similar to the previous test, Test 2.1 is designed to evaluate the value of the weight information. We sampled 20 test functions from the exact same range as the one used for training. We again exclude the condition with novel detection (NAF-w/o-Weight) from this test as there will not be novel functions. The log regret results are presented in Figure 9, left panel. NAF+ and NAF-w/o-Weight start from an overall better performance than the others. Especially, NAF+ converges to optimal performance around the second iteration, and NAF-w/o-Weight converges around the fifth iteration, showing NAF+ still benefits from the additional weight information as a part of the input. We observed that TAF with 15 prior models struggled to improve significantly in this case. This might be because of the higher number of parameters (2 in Test 1 and 6 in Test 2). To cover this number of dimensions, potentially more models will be required. Furthermore, same as the previous test, Standard BO starts from a worse performance and gradually converges toward optimal performance. Finally, to further analyze the benefit of incorporating the weight information, we run independent-samples t-tests to specifically compare the resulting performance of NAF+ and NAF-w/o-Weight at all iterations. We found significant differences in iterations 1-4, all $p<0.05$. This shows that the weight information supports more efficient adaptation in the earlier iterations.
@@
Test 2.2 evaluates the value of merging EI and the novelty detector. To create a scenario where the new user is completely outside of the training range, we significantly shift the range of sampling the model parameters. The mean for sampling synthetic users are set as $[180,120,0.0145,1.6,0.03,2.0]$ for $a,b,\alpha_{x},\sigma_{\alpha_{x}},\alpha_{y},\sigma_{\alpha_{y}}]$, respectively, and the standard deviation remains the same. These ranges are significantly different from the parameter settings for training, ensuring all sampled users are novel. We sampled 20 test functions from the test range. The log regret results are presented in Figure 9, right panel. We find that NAF+ and NAF-w/o-EI start from overall better performance, but NAF-w/o-EI does not improve significantly throughout the iterations, while NAF+ continually improves. This, again, highlights the potential issue of not incorporating an additional EI to handle out-of-distribution cases: naf can not generalize to other cases once it is fully trained to a certain range. TAF and standard BO perform well in this task. However, they start from a worse performance. BO, especially, requires 4 to 5 iterations to catch up with other conditions.
@@ C.5. Summary of Test 2
In summary, both Test 1 (a controlled synthetic benchmark) and Test 2 (a more realistic simulation grounded in HCI literature) converge on the same conclusion: NAF+, through its use of explicit weight input and novelty-aware EI, achieves greater sample efficiency and robustness across a wide range of user profiles. The positive outcomes of Test 2 further motivate our user study, which targets typing performance on mid-air keyboards, highly relevant to Test 2 – a touch-based keyboard for mobile devices.
Figure 9. The results of Test 2 (simulating softkeyboard typing): Test 2.1 (evaluating weight information) is shown in left, and Test 2.2 (evaluating expected improvement) is shown in right. The value plot here is the $log_{10}$ of the regret. In Test 2.1, the * signs denote the iterations (1-4) where a significant difference between NAF+ and NAF-w/o-Weight was found. The results show that NAF+ benefits from the weight information and thus can achieve better performance than NAF-w/o-Weight in earlier iterations. On the other hand, NAF+ can achieve better results compared to NAF-w/o-EI when facing novel tasks, particularly in later iterations.